OpenStudy (anonymous):

magnetic filed at the centre of regular polygons which are formed by constant current carrying wire in equilateral triangle

OpenStudy (anonymous):



OpenStudy (anonymous):

1)\[\mu oI/2\sqrt{3} \pi a\] 2)\[27\mu oI / 2\pi a\] 3)\[9\mu oI/2\pi a\] 4)\[\sqrt{3} \mu oI/\pi a\]

OpenStudy (anonymous):

option 3 i get

OpenStudy (anonymous):